OpenStudy (anonymous):

Solve. 7x2 – 35x = 168

OpenStudy (lgbasallote):

\[\LARGE \mathtt{\text{divide each term by 7 :D}}\]

OpenStudy (anonymous):

7x2-35x=168 x2-5x=24 x2-5x-24=0 x=-3

OpenStudy (anonymous):

x^2-5x-24=0 Factor to (x-8)(x+3)=0 --- x=8 or -3